Write the Equation of the line and graph:
m= -2/3, through the point (-2, -4)

Answer by checkley71(8403)

Y=mX+b
-4=-2/3*-2+b
-4=+4/3+b
b=-4-4/3
b=-16/3
THUS THE EQUATION IS
Y=2/3X-16/3